OpenStudy (anonymous):

Solve using quadratic formula 2x^2-x=-6

OpenStudy (bobbyleary):

\[2x^2-x+6=0\]Quadratic Formula (memorize this):\[x = {-b \pm \sqrt{b^2 - 4ac} \over{2a}}\]In your problem, a = 2, b = -1, and c = 6. Substitute these numbers into the formula.\[x = {-(-1) \pm \sqrt{(-1^2)-4(2)(6)} \over 2(2)}={1 \pm \sqrt{1-48} \over 4} = {1 \over 4} \pm {i \sqrt{47} \over 4}\]